# DisenSemi: Semi-supervised Graph Classification via Disentangled Representation Learning

### v1

@@ V-B1 Overall Comparison
We present the quantitative results of semi-supervised graph classification with $30\%$ label ratio in Table I. According to the results, the following observations can be derived.
@@
  • Our model DisenSemi achieves the best performance on all ten datasets, which significantly demonstrates the effectiveness of our proposed model. From the results, we deem that the improvement over state-of-the-art methods not only from learning disentangled representation of the graph, but also from explicitly regularizing the rationale that fits well between supervised and unsupervised learning models for semi-supervised graph classification.
@@ V-B2 Performance on Different Labeling Ratio
We vary the labeling ratio of training data to show the performance of different models. Here we omit the traditional methods since they do not have competitive performance. Fig. 4 illustrates the performance w.r.t. different labeling ratios of data on four datasets. We find that:
  • The performance of all models generally improves as the number of available labeled data increases. This demonstrates that the labeling ratio of training data is an important factor of the model. Meanwhile, the growth is not linear, which means unsupervised data also plays a key role to boost the model performance.
@@
  • DisenSemi consistently outperforms other baselines over the different labeling ratios of data. Especially when the labeled data is scarce (e.g. less than $70\%$), the superiority of our model indicates that disentangled graph representation with consistency regularization can effectively improve the semi-supervised classification performance.
@@ V-C1 Effect of GNN-based Encoder
In DisenSemi, we encode each factor graph via the proposed GNN with its own edge coefficient. To explore whether DisenSemi can derive benefits from the proposed GNN-based encoder, we compare five types of well-known GNN operators, namely SAGEConv [59], GATConv [60], GINConv [61], GCNConv [62] and GraphConv [49]. Notice that for SAGEConv*, GATConv* and GINConv*, we ignore the edge coefficient and randomly remove $15\%$ edges to obtain different factor graphs. For GraphConv*, we retain the edge coefficients and randomly remove edges separately. The performance of different GNN operators on all ten datasets is shown in Table II. We make the following observations from the results.
@@ V-C3 Effect of Consistency Regularization
Disentangled consistency regularization helps to transfer the learned disentangled graph representations from the unsupervised encoder to the supervised encoder. To further investigate how the consistency regularization facilitates the model performance, we compare DisenSemi with the following two variants:
@@
The result of DisenSemi and its variants are shown in Table III and we have the following observations:
@@ V-D Parameter Sensitivity (RQ3)
We also examine the sensitivity of the proposed DisenSemi to various hyper-parameters.
@@ V-D1 Effect of the Number of Factor Graphs
To analyze whether DisenSemi can benefit from disentangled representation, we study the performance of the model with varying numbers of factor graphs. In particular, we search the number of factor graphs in the range of $\{2^{0},2^{1},\dots,2^{4}\}$. Fig. 5 summarizes the experimental results w.r.t. different numbers of factor graphs on four datasets and we find that:
@@
  • However, when the number of factor graphs is too large (i.e., $K\geq 16$), the model performance will degrade slowly. This might be caused by applying a too complex semantic structure for the graph.
@@ V-D2 Effect of the Number of Message Passing Layers
For each disentangled factor graph, we investigate how the depth of DisenSemi affects the performance. In particular, we stack varying numbers of message passing layers in the range of $[0,5]$. Fig. 6 shows the experimental results w.r.t. different number of message passing layers in four datasets (i.e., MUTAG, PROTEINS, IMDB-BINARY and REDDIT-BINARY) and we have the following findings:
@@
  • More message passing layers will yield influence from high-order neighbors in the graph and increase the depth of the model. Clearly, two layers of a model can achieve better performance than one layer. It illustrates the importance of messages passing between nodes to capture the structure of each factor graph.
@@ V-D3 Effect of Relative Weight between Losses
We investigate the sensitivity of relative weight between losses, namely $\lambda$ and $\gamma$. In particular, we vary the value of $\lambda$ and $\gamma$ from 00 to $1$ on PROTEINS and COLLAB. As shown in Fig. 7, unsupervised module and disentangled consistency regularization play an important role in the objective and there is an evident performance drop when $\lambda,\gamma=0$. As $\lambda$ and $\gamma$ get larger, the performance is relatively stable, and we can get the best results when $\lambda,\gamma=1e-3$.
@@ V-E1 Visualization of Representation Correlation
Besides the quantitative evaluation, we also plot the absolute values of the correlations between the elements of the 128-dimensional graph representations. Fig. 8 shows the correlations analysis of representation obtained from GraphCL, DGCL and our proposed DisenSemi with four factor graphs in the MUTAG and PROTEINS dataset. We find that:
@@
  • The representation generated from DisenSemi presents a more block-wise correlation pattern, indicating that four factor graphs of DisenSemi are likely capturing mutually exclusive information to provide discriminant learned features for the target task.
@@ V-E2 Case Study
For a more intuitive understanding of the disentanglement, we provide examples of the generated factor graphs in DisenSemi. Fig. 9 presents the original graph and those disentangled factor graphs. The edges of the visualized disentangled factor graphs are highlighted with different colors after we set the coefficient threshold in the original graph. For example, in the MUTAG dataset, the task is to predict the molecule’s mutagenicity on Salmonella typhimurium for a collection of nitroaromatic compounds. We can see different parts of a molecule graph playing different roles in the prediction task. This also justifies the reliability of our generated factor graphs in getting the disentangled graph representation.
@@ VI Conclusion
In this paper, we study semi-supervised graph classification, which is a fundamental problem in graph-structured data mining. For transferring suitable knowledge from the unsupervised model to the supervised model, we propose a novel framework termed DisenSemi, which learns disentangled representation to capture multiple graph characteristics stemming from different aspects. Specifically, our DisenSemi consists of a supervised model and an unsupervised model. For both models, we design a disentangled graph encoder to extract factor-wise graph representation and train two models with supervised objective and MI-based constraints, respectively. Then we propose an MI-based disentangled consistency regularization to identify the rationale that aligns well between two models for the current graph classification task and transfer corresponding knowledge semantically. Extensive experiments on ten benchmark graph classification datasets demonstrate the efficacy of our DisenSemi. However, the number of factor graphs varies among datasets and needs to be searched to find the best value. So in future work, we will apply our DisenSemi to more realistic semi-supervised classification scenarios and automatically select factor graph numbers in a bi-level optimization framework.
@@ Acknowledgement
The authors are grateful to the anonymous reviewers for critically reading this article and for giving important suggestions to improve this article.

### v2

@@ V-B1 Overall Comparison
Table I provides the quantitative results of semi-supervised graph classification using a $30\%$ label ratio. From these results, we can draw the following conclusions.
@@
  • Our DisenSemi, outperforms all other models across the ten datasets, showcasing its superior effectiveness. The results indicate that the advancement beyond state-of-the-art methods stems not only from learning disentangled representation of the graph, but also from explicitly regularizing the rationale that fits well between supervised and unsupervised learning models for semi-supervised graph classification.
@@ V-B2 Performance on Different Labeling Ratio
We vary the labeling ratio of training data to compare the performance of different models. Traditional methods are omitted due to their lack of competitive performance. Fig. 3 illustrates the performance w.r.t. different labeling ratios of data on four datasets. We observe that:
  • As the quantity of labeled graph data increases, the performance of all models tends to improve. This demonstrates that the labeling ratio of training data is an important factor of the model. Meanwhile, the growth is not linear, which means unsupervised data also plays a key role to boost the model performance.
@@
  • DisenSemi shows superior performance compared to other baselines for varying labeling ratios of data. Especially when the labeled data is scarce (e.g. less than $70\%$), the superiority of our model indicates that disentangled graph representation with consistency regularization can effectively improve the semi-supervised classification performance.
@@ V-C1 Effect of GNN-based Encoder
In DisenSemi, we encode each factor graph via the proposed GNN with its own edge coefficient. To explore whether DisenSemi can derive benefits from the proposed GNN-based encoder, we compare five types of well-known GNN operators, namely SAGEConv [54], GATConv [55], GINConv [56], GCNConv [57] and GraphConv [44]. Notice that for SAGEConv*, GATConv* and GINConv*, we ignore the edge coefficient and randomly remove $15\%$ edges to obtain different factor graphs. For GraphConv*, we retain the edge coefficients and randomly remove edges separately. The performance of different GNN operators on all ten datasets is shown in Table II. From the results, we can draw the following conclusions.
@@ V-C3 Effect of Consistency Regularization
Disentangled consistency regularization helps to transfer the learned disentangled graph representations from the unsupervised encoder to the supervised encoder. To further investigate how the consistency regularization facilitates the model performance, we evaluate DisenSemi against the following two variants:
@@
Table III displays the outcomes for DisenSemi and its variants. The observations are as follows:
@@ V-D Parameter Sensitivity (RQ3)
We also analyze how the proposed DisenSemi responds to different hyper-parameter settings.
@@ V-D1 Effect of the Number of Factor Graphs
To assess the advantages of disentangled representation for DisenSemi, we evaluate the model’s performance with different quantities of factor graphs. Specifically, we explore the number of factor graphs in the set $\{2^{0},2^{1},\dots,2^{4}\}$. Figure 4 presents a summary of the experimental outcomes across four datasets, revealing the following insights:
@@
  • Nonetheless, if the number of factor graphs is excessively large (e.g., $K\geq 16$), the model’s performance tends to decline gradually. This degradation may result from employing an overly complex semantic structure for the graph.
@@ V-D2 Effect of the Number of Message Passing Layers
To examine the impact of depth on performance for each disentangled factor graph, we test the DisenSemi model with a range of message passing layers from $[0,5]$. Figure 5 illustrates the results of these experiments across four datasets—MUTAG, PROTEINS, IMDB-BINARY, and REDDIT-BINARY. The findings are as follows:
@@
  • Increasing the number of message passing layers incorporates information from higher-order neighbors and deepens the model. Evidently, a model with two layers performs better than one with only one layer. It illustrates the importance of messages passing between nodes to capture the structure of each factor graph.
@@ V-D3 Effect of Relative Weight between Losses
We investigate the sensitivity of relative weight between losses, namely $\lambda$ and $\gamma$. In particular, we adjust the value of $\lambda$ and $\gamma$ from 00 to $1$ on PROTEINS and COLLAB. Figure 6 demonstrates that both the unsupervised module and disentangled consistency regularization play an important role in the objective and there is an evident performance drop when $\lambda,\gamma=0$. As $\lambda$ and $\gamma$ get larger, the performance is relatively stable, and we can get the best results when $\lambda,\gamma=1e-3$.
@@ V-E1 Visualization of Representation Correlation
Besides the quantitative evaluation, we also visualize the absolute correlations between the components of the 128-dimensional graph representations. Fig. 7 shows the correlations analysis of representation obtained from GraphCL, DGCL and our proposed DisenSemi with four factor graphs in the MUTAG and PROTEINS dataset. We find that:
@@
  • The representation produced by DisenSemi demonstrates a distinct block-wise correlation pattern, suggesting that the four factor graphs of DisenSemi are likely capturing mutually exclusive information. This allows for the extraction of discriminative features for the target task.
@@ V-E2 Case Study
To illustrate the disentanglement process more clearly, we present examples of the factor graphs generated by DisenSemi. Figure 8 shows the original graph alongside the disentangled factor graphs. The edges of the visualized disentangled factor graphs are highlighted with different colors after we set the coefficient threshold in the original graph. For example, in the MUTAG dataset, the task is to predict the molecule’s mutagenicity on Salmonella typhimurium for a collection of nitroaromatic compounds. We can see different parts of a molecule graph playing different roles in prediction. This also justifies the reliability of our generated factor graphs in getting the disentangled graph representation.
@@ VI Conclusion
This paper explores the problem of semi-supervised graph classification, a core problem in the field of graph-structured data analysis. For transferring suitable knowledge from the unsupervised model to the supervised model, we propose a novel framework termed DisenSemi, which learns disentangled representation to capture multiple graph characteristics stemming from different aspects. Specifically, our DisenSemi consists of a supervised model and an unsupervised model. For both models, we design a disentangled graph encoder to extract factor-wise graph representation and train two models with supervised objective and MI-based constraints, respectively. Then we propose an MI-based disentangled consistency regularization to identify the rationale that aligns well between two models for the current graph classification task and transfer corresponding knowledge semantically. Extensive experiments on ten benchmark graph classification datasets demonstrate the efficacy of our DisenSemi. However, the number of factor graphs varies among datasets and needs to be searched to find the best value. In future research, we plan to extend our DisenSemi to more complex semi-supervised classification scenarios and automatically select factor graph numbers in a bi-level optimization framework.
@@ Acknowledgement
The paper is supported by “the Fundamental Research Funds for the Central Universities” in UIBE (Grant No. 23QN02) and the National Natural Science Foundation of China (NSFC Grant Numbers 62306014 and 62276002).